OpenStudy (anonymous):

The variables x and y vary directly. Use the given values to write an equation that relates x and y: x = 4, y = -16

OpenStudy (anonymous):

x=ky, where k is a constant. 4=-16k k=-1/4 x=-1/4 y or y =-4x

OpenStudy (the_fizicx99):

I believe the correct formula is y = kx -16 = 4x /4 /4 x = -4